OpenStudy (31356):

How many boxes will be made from a grid with 3x^6 and 5x^4 columns? MEDAL REWARDED!!

OpenStudy (31356):

Any ideas? No answer choices here.

OpenStudy (31356):

@jim_thompson5910 @whpalmer4 @preetha

OpenStudy (whpalmer4):

How many columns does the grid have? Are you missing the word "rows" in the problem statement, perhaps?

OpenStudy (31356):

No

OpenStudy (whpalmer4):

well, a grid only has one number of columns, not two. it also has a number of rows.

OpenStudy (31356):

I think the answer is talking about 3x^6*5x^4 right?

OpenStudy (whpalmer4):



OpenStudy (31356):

Since it has something to do with factoring using distributive property?

OpenStudy (31356):

What are the exponents for then?

OpenStudy (whpalmer4):

that would be my guess, but I'd be happier if the problem was stated in a fashion that made sense. Having two values listed for columns does not.

OpenStudy (31356):

Oh yea 3x^6 rows you are right

OpenStudy (31356):

Mislooked :(

OpenStudy (whpalmer4):

the number of rows is \(3x^6\) and the number of columns is \(5x^4\) Multiply those quantities together to get the number of boxes.

OpenStudy (31356):

15x^10?

OpenStudy (31356):

Can you help me with another one?

OpenStudy (whpalmer4):

\[3x^6*5x^4 = 3*5*x^{6+4} = 15x^{10}\checkmark\]

OpenStudy (whpalmer4):

Sure, why not?

OpenStudy (31356):

This one's hard, in my opinion.

OpenStudy (whpalmer4):

What are you supposed to do?

OpenStudy (31356):

Simplify the following expression

OpenStudy (whpalmer4):

Okay, go through and apply the distributive property to get rid of the parentheses. what do you get?

OpenStudy (31356):

I am not sure how to do it honestly.

OpenStudy (whpalmer4):

Let's walk through it step by step: \[5h^4(5+h^2)+h^3(3h^3+6h)-(h^2+4)\] \[5h^4(5+h^2) = \]

OpenStudy (31356):

Is it 5h^4(h^2+5) I am not sure @whpalmer4

OpenStudy (whpalmer4):

Use the distributive property: \[a(b+c) = a*b + a*c\]

OpenStudy (31356):

Okay

OpenStudy (31356):

25h^6?

OpenStudy (31356):

I am not sure @whpalmer4

OpenStudy (whpalmer4):

Okay, compare the expressions: \[a(b+c) = a*b + a*c\]\[5h^4(5+h^2) = \] What is \(a=\)? What is \(b=\)? What is \(c=\)?

OpenStudy (31356):

I don't know how can you multiply 5h^4 by 5?

OpenStudy (31356):

Or would it be 25h^4+6h^6? @whpalmer4

OpenStudy (whpalmer4):

\[5h^4*5 = 5*5h^4=25h^4\]

OpenStudy (31356):

Okay, I was correct , so as the other one?

OpenStudy (whpalmer4):

\[5h^4*h^2 = 5h^{4+2} = 5h^6\]

OpenStudy (31356):

Oh okay

OpenStudy (31356):

What do we do next?

OpenStudy (whpalmer4):

So, continue through the expression: \[h^3(3h^3+6h) =\]

OpenStudy (31356):

Okay

OpenStudy (31356):

3h^6+6h^4? @whpalmer4

OpenStudy (whpalmer4):

Very good. \[-(h^2+4) =\]This is slightly tricky!

OpenStudy (31356):

How are we suppose to do this?

OpenStudy (whpalmer4):

here's a hint: \[-(h^2+4) = -1(h^2+4)\]

OpenStudy (31356):

Oh yeah!

OpenStudy (31356):

-1h^2+3?

OpenStudy (31356):

@whpalmer4

OpenStudy (whpalmer4):

where'd that 3 come from?

OpenStudy (31356):

-1+4?

OpenStudy (whpalmer4):

\[-1(h^2+4)\]\[a(b+c) = a*b + a*c\]want to try again?

OpenStudy (31356):

Sure

OpenStudy (31356):

-1h^2-4? I don't know Sorry, lost connection

OpenStudy (whpalmer4):

\[-1(h^2+4) = -1*h^2 -1*4 = -h^2-4\]

OpenStudy (31356):

Oh okay, so close

OpenStudy (whpalmer4):

Okay, so now we need to collect all the pieces and stick it together: \[5h^4(5+h^2)+h^3(3h^3+6h)-(h^2+4) = 25h^4+5h^6+3h^6+6h^4-h^2-4\] Can you collect any like terms? You were correct by the way, didn't mean to suggest otherwise!

OpenStudy (31356):

Then do we just add like terms?

OpenStudy (whpalmer4):

\[-1h^2 = -h^2\]

OpenStudy (whpalmer4):

yes, that's what you do...

OpenStudy (31356):

Okay

OpenStudy (whpalmer4):

and then write them in order of descending exponents...

OpenStudy (31356):

:D

OpenStudy (31356):

Is it 8h^12+31h^10-h^2-4? @whpalmer4

OpenStudy (whpalmer4):

No, how did you get \(h^{12}\) and \(h^{10}\) from addition where none of the terms had those powers?

OpenStudy (31356):

I don't know I think I mixed it up somewhere.

OpenStudy (31356):

So what am I suppose to do then?

OpenStudy (whpalmer4):

you multiplied terms instead of adding them... \[3x^2+5x^2 = (3+5)x^2 = 8x^2\]

OpenStudy (31356):

Oh okay

OpenStudy (whpalmer4):

you need to review algebraic manipulation of things with exponents, I think :-)

OpenStudy (31356):

I was absent on that day

OpenStudy (whpalmer4):

it happens. but you still need to know the material

OpenStudy (31356):

So is it the same thing just both of the h's are 24?

OpenStudy (whpalmer4):

write out the answer

OpenStudy (31356):

8h^24+31h^24-h^2-4?

OpenStudy (31356):

@whpalmer4

OpenStudy (whpalmer4):

How are you getting these enormous exponents?!? You don't change the exponent when adding or subtracting!

OpenStudy (31356):

Oh really so it's the same thing ......

OpenStudy (whpalmer4):

Look, adding \[3x^2 + 2x^2 = 3*x*x + 2*x*x\]That doesn't give you \[6*x*x*x*x!\]

OpenStudy (31356):

Okay

OpenStudy (whpalmer4):

it gives you \[(x*x + x*x + x*x) + (x*x+ x*x) = 5*x*x = 5x^2\]

OpenStudy (31356):

I am just so bad at this unit!

OpenStudy (whpalmer4):

You need to talk to your teacher about getting the material you missed.

OpenStudy (31356):

I know planning tomorrow

OpenStudy (whpalmer4):

This is absolutely crucial bedrock algebra, and you'll be in deep doo-doo if you don't understand it...

OpenStudy (31356):

That's why I need your help due to test tomorrow

OpenStudy (31356):

So we just leave the exponents as usual right?

OpenStudy (31356):

Or the same thing when it once was?